OpenStudy (anonymous):

The limit represents the derivative of some function f at some number a. State such an f and a. lim-->0 (cos(π + h) + 1)/h f(x) = cos(x), a = π f(x) = cos(x), a = 0 f(x) = cos(x), a = π/4 f(x) = cos(x), a = π/3 f(x) = cos(x), a = π/6

OpenStudy (aum):

\[ f(x) = \cos(x) \\ f'(a) = \lim_{h \rightarrow 0}\frac{\cos(a+h)-\cos(a)}{h} \\ \text{For } a = \pi \\ f'(\pi) = \lim_{h \rightarrow 0}\frac{\cos(\pi+h)-\cos(\pi)}{h} = \lim_{h \rightarrow 0}\frac{\cos(\pi+h)+1}{h} \]